# RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning

### v1

@@ Abstract
We introduce a novel Inverse Reinforcement Learning (IRL) approach that overcomes limitations of fixed reward assignments and constrained flexibility in implicit reward regularization. By extending the Maximum Entropy IRL framework with a squared temporal-difference (TD) regularizer and adaptive targets, dynamically adjusted during training, our method indirectly optimizes a reward function while incorporating reinforcement learning principles. Furthermore, we integrate distributional RL to capture richer return information. Our approach achieves state-of-the-art performance on challenging MuJoCo tasks, demonstrating expert-level results on the Humanoid task with only 3 demonstrations. Extensive experiments and ablation studies validate the effectiveness of our method, providing insights into adaptive targets and reward dynamics in imitation learning. [^1] The code is available at https://github.com/adibka/RIZE.
@@ 1   Introduction
Inverse Reinforcement Learning (IRL) [Abbeel and Ng, 2004] is a foundational paradigm in artificial intelligence, enabling agents to acquire complex behaviors by observing expert demonstrations. This approach has catalyzed progress in diverse domains such as robotics [Osa et al., 2018], autonomous driving [Knox et al., 2023], and drug discovery [Ai et al., 2024]. By eliminating the need for explicitly defined reward functions, IRL provides a practical framework for training agents in environments where designing such functions is infeasible.
@@
A prominent framework in IRL is Maximum Entropy (MaxEnt) IRL [Ziebart, 2010], which underpins many state-of-the-art (SOTA) IL methods. Prior works have combined MaxEnt IRL with adversarial training [Ho and Ermon, 2016, Fu et al., 2018] to minimize divergences between agent and expert distributions. However, these adversarial methods often suffer from instability during training. To address this, recent research has introduced implicit reward regularization, which indirectly represents rewards via Q-value functions by inverting the Bellman equation. For instance, IQ-Learn [Garg et al., 2021] unifies reward and policy representations using Q-functions with an $L_{2}$-norm regularization on rewards, while LSIQ [Al-Hafez et al., 2023] minimizes the chi-squared divergence between expert and mixture distributions, resulting in a squared temporal difference (TD) error objective analogous to SQIL [Reddy et al., 2020]. Despite its effectiveness, this method has limitations: LSIQ assigns fixed rewards (e.g., +1 for expert samples and -1 for agent samples), which constrains flexibility by treating all tasks and state-action pairs uniformly, limiting performance and requiring additional gradient steps for convergence.
We present an extension of implicit reward regularization under the MaxEnt IRL framework through two methodological innovations: (1) adaptive target rewards ($\lambda^{\pi_{E}},\lambda^{\pi}$) that dynamically adjust during training to replace static targets, enabling context-sensitive alignment of expert and policy data, and (2) distributional RL integration, where value distributions $Z^{\pi}(s,a)$ [Bellemare et al., 2017] are trained to capture richer return information while preserving theoretical consistency with Q-values, as the expectation of the value distribution is used for policy and value function optimization. By unifying adaptive target regularization with distributional insights [Zhou et al., 2023], our framework addresses prior rigidity in reward learning and empirically outperforms online IL baselines on MuJoCo benchmarks [Todorov et al., 2012], marking a novel integration of distributional RL into non-adversarial IRL.
Our contributions are as follow:
  • We introduce adaptive target rewards for implicit reward regularization to enable dynamic adjustments during training.
  • We leverage value distributions instead of classical Q-functions to capture richer insights from return distributions.
  • We demonstrate the effectiveness of our approach through extensive experiments on MuJoCo tasks, providing detailed analyses of implicit rewards and adaptive target mechanisms.
2   Related Works
@@
Imitation learning (IL) and inverse reinforcement learning (IRL) [Watson et al., 2023] are foundational paradigms for training agents to mimic expert behavior from demonstrations. Behavioral Cloning (BC) [Pomerleau, 1991], the simplest IL approach, treats imitation as a supervised learning problem by directly mapping states to expert actions. While computationally efficient, BC is prone to compounding errors [Ross and Bagnell, 2011] due to covariate shift during deployment. The Maximum Entropy IRL framework [Ziebart, 2010] addresses this limitation by probabilistically modeling expert behavior as reward maximization under an entropy regularization constraint, establishing a theoretical foundation for modern IRL methods.
@@
The advent of adversarial training marked a pivotal shift in IL methodologies. Ho and Ermon [2016] introduced Generative Adversarial Imitation Learning (GAIL), which formulates imitation learning as a generative adversarial game [Goodfellow et al., 2014] where an agent learns a policy indistinguishable from the expert’s by minimizing the Jensen-Shannon divergence between their state-action distributions. This framework was generalized by Ghasemipour et al. [2019] in f-GAIL, which replaces the Jensen-Shannon divergence with arbitrary $f$-divergences to broaden applicability. Concurrently, Kostrikov et al. [2019] proposed Discriminator Actor-Critic (DAC), improving sample efficiency via off-policy updates and terminal-state reward modeling while mitigating reward bias.
@@
Recent advances have shifted toward methods that bypass explicit reward function estimation. Kostrikov et al. [2020] introduced ValueDICE, an offline IL method that leverages an inverse Bellman operator to avoid adversarial optimization. Similarly, Garg et al. [2021] developed IQ-Learn, which circumvents the challenges of MaxEnt IRL by optimizing implicit rewards derived directly from expert Q-values. A parallel research direction simplifies reward engineering by assigning fixed rewards to expert and agent samples. Reddy et al. [2020] pioneered this approach with Soft Q Imitation Learning (SQIL), which assigns binary rewards to transitions from expert and agent trajectories. Most recently, Al-Hafez et al. [2023] proposed Least Squares Inverse Q-Learning (LSIQ), enhancing regularization by minimizing chi-squared divergence between expert and mixture distributions while explicitly managing absorbing states through critic regularization. Our work builds on these foundations by integrating distributional RL [Dabney et al., 2018a] and introducing a squared TD regularizer with adaptive targets.
@@ 3.1   Preliminary
We consider a Markov Decision Process (MDP) [Puterman, 2014] to model policy learning in Reinforcement Learning (RL). The MDP framework is defined by the tuple $\langle\mathcal{S},\mathcal{A},p_{0},P,R,\gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p_{0}$ the initial state distribution, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ the transition kernel with $P(\cdot\mid s,a)$ specifying the likelihood of transitioning from state $s$ given action $a$, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function, and $\gamma\in[0,1]$ the discount factor which tempers future rewards. A stationary policy $\pi\in\Pi$ is characterized as a mapping from states $s\in\mathcal{S}$ to distributions over actions $a\in\mathcal{A}$. The primary objective in RL [Sutton and Barto, 2018] is to maximize the expected sum of discounted rewards, expressed as $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$. Furthermore, the occupancy measure $\rho_{\pi}(s,a)$ for a policy $\pi\in\Pi$ is given by $(1-\gamma)\pi(a\mid s)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s\mid\pi)$. The corresponding measure for an expert policy, $\pi_{E}$, is similarly denoted by $\rho_{E}$. In Imitation Learning (IL), it is posited that the expert policy $\pi_{E}$ remains unknown, and access is restricted to a finite collection of expert demonstrations, represented as $(s,a,s^{\prime})$.
@@ 3.2   Distributional Reinforcement Learning
Maximum Entropy (MaxEnt) RL [Haarnoja et al., 2018] focuses on addressing the stochastic nature of action selection by maximizing the entropy of the policy, while Distributional RL [Bellemare et al., 2017] emphasizes capturing the inherent randomness in returns. Combining these perspectives, the distributional soft value function $Z^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{Z}$ [Ma et al., 2020] for a policy $\pi\in\Pi$ encapsulates uncertainty in both rewards and actions. It is formally defined as:
$Z^{\pi}(s,a)=\sum_{t=0}^{\infty}\gamma^{t}[R(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot\mid s_{t}))],$ (1)
where $\mathcal{Z}$ represents the space of action-value distributions, $\mathcal{H}(\pi)=\mathbb{E}_{\pi}[-\log\pi(a\mid s)]$ denotes the entropy of the policy, and $\alpha>0$ balances entropy with reward.
[…]
A practical approach to approximating the value distribution $Z$ involves modeling its quantile function $F^{-1}_{Z}(\tau)$, evaluated at specific quantile levels $\tau\in[0,1]$ [Dabney et al., 2018b]. The quantile function is defined as $F^{-1}_{Z}(\tau)=\inf\{z\in\mathbb{R}:\tau\leq F_{Z}(z)\}$, where $F_{Z}(z)=\mathbb{P}(Z\leq z)$ is the cumulative distribution function of $Z$. For simplicity, we denote the quantile-based representation as $Z_{\tau}^{\pi}(s,a):=F^{-1}_{Z}(\tau)$. To discretize this representation, we define a sequence of quantile levels, denoted as $\{\tau_{i}\}_{i=0,...,N-1}$, where $10=\tau_{0}<...<\tau_{N-1}=$ 1. These quantiles partition the unit interval into $N$ fractions. Sampling uniformly from this interval $\tau\sim U(0,1)$, we obtain an approximation of the value distribution $Z^{\pi}_{\tau}(s,a)$, which provides quantile values at these specified levels.
@@ 3.3   Inverse Reinforcement Learning
Given expert trajectory data, Maximum Entropy (MaxEnt) Inverse RL [Ziebart, 2010] aims to infer a reward function $R(s,a)$ from the family $\mathcal{R}=\mathbb{R}^{S\times A}$. Instead of assuming a deterministic expert policy, this method optimizes for stochastic policies $\pi\in\Pi$ that maximize $R$ while matching expert behavior. Ho and Ermon [2016] extend this framework by introducing a convex reward regularizer $\psi:\mathbb{R}^{S\times A}\rightarrow\bar{\mathbb{R}}$, leading to the adversarial objective:
$\max_{R\in\mathcal{R}}\min_{\pi\in\Pi}L(\pi,R)=\mathbb{E}_{\rho_{E}}[R(s,a)]-\mathbb{E}_{\rho_{\pi}}[R(s,a)]-\mathcal{H}(\pi)-\psi(R)$ (2)
IQ-Learn [Garg et al., 2021] departs from adversarial training by implicitly representing rewards through Q-functions $Q\in\Omega$ [Piot et al., 2014]. It leverages the inverse soft Bellman operator $\mathcal{T}^{\pi}$, defined as:
$(\mathcal{T}^{\pi}Q)(s,a)=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a),a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})]$ (3)
For a fixed policy $\pi$, $\mathcal{T}^{\pi}$ is bijective, ensuring a one-to-one correspondence between $Q$-values and rewards: $\mathcal{T}^{\pi}Q=R$ and $Q=(\mathcal{T}^{\pi})^{-1}R$. This allows reframing the MaxEnt IRL objective in Q-policy space as $\max_{Q\in\Omega}\min_{\pi\in\Pi}\mathcal{J}(\pi,Q)$. IQ-Learn simplifies the problem by defining the implicit reward $R_{Q}(s,a)=\mathcal{T}^{\pi}Q(s,a)$ and applying an L2 regularizer $\psi(R_{Q})$. The final objective becomes:
$\mathcal{J}(\pi,Q)$ $=\mathbb{E}_{\rho_{E}}[R_{Q}(s,a)]-\mathbb{E}_{\rho_{\pi}}[R_{Q}(s,a)]-\mathcal{H}(\pi)-c\left(\mathbb{E}_{\rho_{E}}[R_{Q}^{2}]+\mathbb{E}_{\rho_{\pi}}[R_{Q}^{2}]\right).$ (4)

### v2

@@ Abstract
We propose a novel Inverse Reinforcement Learning (IRL) method that mitigates the rigidity of fixed reward structures and the limited flexibility of implicit reward regularization. Building on the Maximum Entropy IRL framework, our approach incorporates a squared temporal-difference (TD) regularizer with adaptive targets that evolve dynamically during training, thereby imposing adaptive bounds on recovered rewards and promoting robust decision-making. To capture richer return information, we integrate distributional RL into the learning process. Empirically, our method achieves expert-level performance on complex MuJoCo tasks, surpassing baseline methods on the Humanoid task with 3 demonstrations. Extensive experiments and ablation studies further validate the effectiveness of the approach and provide insights into reward dynamics in imitation learning.  [^1] Our code is available at https://github.com/adibka/RIZE.
@@ 1 Introduction
Designing effective reward functions remains a fundamental challenge in Reinforcement Learning (RL). While dense, well-shaped rewards can facilitate learning, they often require laborious task-specific tuning and domain expertise, limiting their scalability (Amodei et al., 2016; Peng et al., 2020). To circumvent these limitations, researchers have explored alternative approaches, including sparse rewards upon task completion (Silver et al., 2016), learning from expert trajectories (Schaal, 1996; Ng & Russell, 2000; Osa et al., 2018), human preference-based reward modeling (Christiano et al., 2017; Lee et al., 2021; Hejna & Sadigh, 2023), and intrinsically motivated RL (Oudeyer et al., 2007; Schmidhuber, 2010; Colas et al., 2022). Among these, Inverse Reinforcement Learning (IRL) (Abbeel & Ng, 2004) offers a compelling alternative by inferring reward functions directly from expert demonstrations, bypassing manual reward engineering. IRL has driven breakthroughs in robotics (Osa et al., 2018), autonomous driving (Knox et al., 2023), and drug discovery (Ai et al., 2024).
@@
A prominent framework in IRL is Maximum Entropy (MaxEnt) IRL (Ziebart, 2010), which underpins many state-of-the-art (SOTA) IL methods. Prior works have combined MaxEnt IRL with adversarial training (Ho & Ermon, 2016; Fu et al., 2018) to minimize divergences between agent and expert distributions. However, these adversarial methods often suffer from instability during training. To address this, recent research has introduced implicit reward regularization, which indirectly represents rewards via Q-values by inverting the Bellman equation. For instance, IQ-Learn (Garg et al., 2021) unifies reward and policy representations using Q-functions with an $L_{2}$-norm regularization on rewards, while LSIQ (Al-Hafez et al., 2023) minimizes the chi-squared divergence between expert and mixture distributions, resulting in a squared temporal difference (TD) error objective analogous to SQIL (Reddy et al., 2020). Despite its effectiveness, this method has limitations: LSIQ assigns fixed targets for implicit rewards (e.g., +1 for expert samples and -1 for agent samples), which constrains flexibility by treating all tasks and state-action pairs uniformly, limiting performance and requiring additional gradient steps for convergence.
We propose an extension of implicit reward regularization under the MaxEnt IRL framework, introducing two key advancements. Adaptive Reward Targets: We enhance prior TD-error regularization by introducing learnable targets $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ that dynamically adjust during training. These targets replace static constraints with context-sensitive reward alignment, preventing rewards from over-increasing or over-decreasing. Crucially, our theoretical analysis reveals that these adaptive bounds constrain implicit rewards to well-defined ranges $\left[-\frac{1}{2c}+\min\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\},\;\frac{1}{2c}+\max\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}\right]$, where $c$ is the regularization coefficient. Since implicit rewards derive from Q-values, this regularization indirectly stabilizes policy training. Distributional RL Integration: We incorporate value distributions $Z^{\pi}(s,a)$ (Bellemare et al., 2017) to capture richer uncertainty information in returns, while using their expectations for policy optimization. Though distributional RL has shown success in adversarial IRL (Zhou et al., 2023), it remains unexplored in non-adversarial settings. Our work bridges this gap, demonstrating its efficacy in MaxEnt IRL while preserving theoretical guarantees. Unifying these advances, our framework outperforms online IL baselines on MuJoCo benchmarks (Todorov et al., 2012). Notably, our approach proves vital for complex tasks like Humanoid, where both adaptive rewards and distributional learning are essential.
Our contributions are threefold. First, we introduce adaptive targets for implicit reward regularization, enabling dynamic reward bounds that enhance stability and alignment during training. Second, we integrate value distributions into implicit reward frameworks, capturing richer uncertainty information while preserving theoretical consistency. Third, we empirically validate our approach through extensive MuJoCo experiments, demonstrating superior performance in complex environments.
2 Related Work
@@
Imitation learning (IL) and inverse reinforcement learning (IRL) (Watson et al., 2023) are foundational paradigms for training agents to mimic expert behavior from demonstrations. Behavioral Cloning (BC) (Pomerleau, 1991), the simplest IL approach, treats imitation as a supervised learning problem by directly mapping states to expert actions. While computationally efficient, BC is prone to compounding errors (Ross & Bagnell, 2011) due to covariate shift during deployment. The Maximum Entropy IRL framework (Ziebart, 2010) addresses this limitation by probabilistically modeling expert behavior as reward maximization under an entropy regularization constraint, establishing a theoretical foundation for modern IRL methods.
@@
The advent of adversarial training marked a pivotal shift in IL methodologies. Ho & Ermon (2016) introduced Generative Adversarial Imitation Learning (GAIL), which formulates imitation learning as a generative adversarial game (Goodfellow et al., 2014) where an agent learns a policy indistinguishable from the expert’s by minimizing the Jensen-Shannon divergence between their state-action distributions. This framework was generalized by Ghasemipour et al. (2019) in f-GAIL, which replaces the Jensen-Shannon divergence with arbitrary $f$-divergences to broaden applicability. Concurrently, Kostrikov et al. (2019) proposed Discriminator Actor-Critic (DAC), improving sample efficiency via off-policy updates and terminal-state reward modeling while mitigating reward bias.
@@
Recent advances have shifted toward methods that bypass explicit reward function estimation. Kostrikov et al. (2020) introduced ValueDICE, an offline IL method that leverages an inverse Bellman operator to avoid adversarial optimization. Similarly, Garg et al. (2021) developed IQ-Learn, which circumvents the challenges of MaxEnt IRL by optimizing implicit rewards derived directly from expert Q-values. A parallel research direction simplifies reward engineering by assigning fixed rewards to expert and agent samples. Reddy et al. (2020) pioneered this approach with Soft Q Imitation Learning (SQIL), which assigns binary rewards to transitions from expert and agent trajectories. Most recently, Al-Hafez et al. (2023) proposed Least Squares Inverse Q-Learning (LSIQ), enhancing regularization by minimizing chi-squared divergence between expert and mixture distributions while explicitly managing absorbing states through critic regularization. Our work builds on these foundations by integrating distributional RL (Dabney et al., 2018a) and introducing a squared TD regularizer with adaptive targets.
@@ 3.1 Preliminary
We consider a Markov Decision Process (MDP) (Puterman, 2014) to model policy learning in Reinforcement Learning (RL). The MDP framework is defined by the tuple $\langle\mathcal{S},\mathcal{A},p_{0},P,R,\gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p_{0}$ the initial state distribution, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ the transition kernel with $P(\cdot\mid s,a)$ specifying the likelihood of transitioning from state $s$ given action $a$, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function, and $\gamma\in[0,1]$ the discount factor which tempers future rewards. A stationary policy $\pi\in\Pi$ is characterized as a mapping from states $s\in\mathcal{S}$ to distributions over actions $a\in\mathcal{A}$. The primary objective in RL (Sutton & Barto, 2018) is to maximize the expected sum of discounted rewards, expressed as $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$. Furthermore, the occupancy measure $\rho_{\pi}(s,a)$ for a policy $\pi\in\Pi$ is given by $(1-\gamma)\pi(a\mid s)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s\mid\pi)$. The corresponding measure for an expert policy, $\pi_{E}$, is similarly denoted by $\rho_{E}$. In Imitation Learning (IL), the expert policy $\pi_{E}$ is typically unknown, and only a finite set of expert demonstrations is available, rather than explicit reward feedback from the environment.
@@ 3.2 Distributional Reinforcement Learning
Maximum Entropy (MaxEnt) RL (Haarnoja et al., 2018) focuses on addressing the stochastic nature of action selection by maximizing the entropy of the policy, while Distributional RL (Bellemare et al., 2017) emphasizes capturing the inherent randomness in returns. Combining these perspectives, the distributional soft value function $Z:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{Z}$ (Ma et al., 2020) for a policy $\pi\in\Pi$ encapsulates uncertainty in both rewards and actions, with $\mathcal{Z}$ representing the space of action-value distributions. It is formally defined as:
$Z(s,a)=\sum_{t=0}^{\infty}\gamma^{t}[R(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot\mid s_{t}))],$ (1)
where $\mathcal{H}(\pi)=\mathbb{E}_{\pi}[-\log\pi(a\mid s)]$ denotes the entropy of the policy, and $\alpha>0$ balances entropy with reward.
The distributional soft Bellman operator $\mathcal{B}_{D}^{\pi}:\mathcal{Z}\rightarrow\mathcal{Z}$ for a given policy $\pi$ is introduced as $(\mathcal{B}_{D}^{\pi}Z)(s,a)\overset{D}{=}R(s,a)+\gamma[Z(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}\mid s^{\prime})]$, where $s^{\prime}\sim P(\cdot\mid s,a)$, $a^{\prime}\sim\pi(\cdot\mid s^{\prime})$, and $\overset{D}{=}$ signifies equality in distribution. Notably, this operator exhibits contraction properties under the p-Wasserstein metric, ensuring convergence to the optimal distributional value function.
A practical approach to approximating the value distribution $Z$ involves modeling its quantile function $F^{-1}_{Z}(\tau)$, evaluated at specific quantile levels $\tau\in[0,1]$ (Dabney et al., 2018b). The quantile function is defined as $F^{-1}_{Z}(\tau)=\inf\{z\in\mathbb{R}:\tau\leq F_{Z}(z)\}$, where $F_{Z}(z)=\mathbb{P}(Z\leq z)$ is the cumulative distribution function of $Z$. For simplicity, we denote the quantile-based representation as $Z_{\tau}(s,a):=F^{-1}_{Z}(\tau)$. To discretize this representation, we define a sequence of quantile levels, denoted as $\{\tau_{i}\}_{i=0,...,N-1}$, where $0=\tau_{0}<...<\tau_{N-1}=1$. These quantiles partition the unit interval into $N$ fractions. Sampling uniformly from this interval $\tau\sim U(0,1)$, we obtain an approximation of the value distribution $Z_{\tau}(s,a)$, which provides quantile values at these specified levels.
@@ 3.3 Inverse Reinforcement Learning
Given expert trajectory data, Maximum Entropy (MaxEnt) Inverse RL (Ziebart, 2010) aims to infer a reward function $R(s,a)$ from the family $\mathcal{R}=\mathbb{R}^{S\times A}$. Instead of assuming a deterministic expert policy, this method optimizes for stochastic policies $\pi\in\Pi$ that maximize $R$ while matching expert behavior. GAIL (Ho & Ermon, 2016) extends this framework by introducing a convex reward regularizer $\psi:\mathbb{R}^{S\times A}\rightarrow\bar{\mathbb{R}}$, leading to the adversarial objective:
$\max_{R\in\mathcal{R}}\min_{\pi\in\Pi}L(\pi,R)=\mathbb{E}_{\rho_{E}}[R(s,a)]-\mathbb{E}_{\rho_{\pi}}[R(s,a)]-\mathcal{H}(\pi)-\psi(R)\,.$ (2)
IQ-Learn (Garg et al., 2021) departs from adversarial training by implicitly representing rewards through Q-functions $Q\in\Omega$ (Piot et al., 2014). It leverages the inverse soft Bellman operator $\mathcal{T}^{\pi}$, defined as:
$(\mathcal{T}^{\pi}Q)(s,a)$ $=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a),a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})].$ (3)
For a fixed policy $\pi$, $\mathcal{T}^{\pi}$ is bijective, ensuring a one-to-one correspondence between $Q$-values and rewards: $\mathcal{T}^{\pi}Q=R$ and $Q=(\mathcal{T}^{\pi})^{-1}R$. This allows reframing the MaxEnt IRL objective (2) in Q-policy space as $\max_{Q\in\Omega}\min_{\pi\in\Pi}\mathcal{J}(\pi,Q)$. IQ-Learn simplifies the problem by defining the implicit reward $R_{Q}(s,a)=\mathcal{T}^{\pi}Q(s,a)$ and applying an L2 regularizer $\psi(R_{Q})$. The final objective becomes:
$\max_{Q\in\Omega}\min_{\pi\in\Pi}\mathcal{J}(\pi,Q)$ $=\;\mathbb{E}_{\rho_{E}}[R_{Q}(s,a)]-\mathbb{E}_{\rho_{\pi}}[R_{Q}(s,a)]-\alpha\mathcal{H}(\pi)$ (4)
$\qquad-c\Big{[}\mathbb{E}_{\rho_{E}}[R_{Q}(s,a)^{2}]+\mathbb{E}_{\rho_{\pi}}[R_{Q}(s,a)^{2}]\Big{]}\,.$